# A Differentiable Model for Optimizing the Genetic Drivers of Synaptogenesis

### v2

@@ Abstract
There is a growing consensus among neuroscientists that many neural circuits critical for survival result from a process of genomic decompression, hence are constructed based on the information contained within the genome. Aligning with this perspective, we introduce SynaptoGen, a novel computational framework designed to bring the advent of synthetic biological intelligence closer, facilitating the development of neural biological agents through the precise control of genetic factors governing synaptogenesis. SynaptoGen represents the first model in the well-established family of Connectome Models (CMs) to offer a possible mechanistic explanation of synaptic multiplicity based on genetic expression and protein interaction probabilities, modeling connectivity with unprecedented granularity. Furthermore, SynaptoGen connects these genetic factors through a differentiable function, effectively working as a neural network in which each synaptic weight is computed as the average number of synapses between neurons, multiplied by its corresponding conductance, and derived from a specific genetic profile. Differentiability is a critical feature of the framework, enabling its integration with gradient-based optimization techniques. This allows SynaptoGen to generate patterns of genetic expression and/or genetic rules capable of producing pre-wired biological agents tailored to specific tasks. The framework is validated in simulated synaptogenesis scenarios with varying degrees of biological plausibility. It successfully produces biological agents capable of solving tasks in four different reinforcement learning benchmarks, consistently outperforming the state-of-the-art and a control baseline designed to represent populations of neurons where synapses form freely, i.e., without guided manipulations. Despite its novel contributions, SynaptoGen exhibits gaps in biological realism that must be addressed before it can be applied in real-world contexts. For this reason, we discuss strategies to mitigate these limitations. Still, SynaptoGen has the potential to accelerate research in synthetic biological intelligence by advancing our understanding of its mechanisms and enabling the development of novel computational platforms.
@@ 1 Introduction
Let us consider a scenario in which, during brain development, h1. it becomes possible to manipulate, prior to the formation of synapses, the gene expression profiles of individual neurons. Furthermore, let us assume that h2. we possess the capability to influence these expression profiles in a manner that directs synaptogenesis towards a specific neuronal network topology. In this context, h3. we might be able to obtain the optimal computational graph, expressed as a composition of functions that represent the behaviour of neurons, required to solve a task of interest. Small living organisms, or organoids (Bhaduri et al., 2020), could be, in principle, genetically programmed to develop into neuronal networks capable of solving pre-specified tasks. Such technology would lead to disruptive applications–e.g., extreme low-power computing, micro-devices for the control of biological systems, or novel biological testing platforms capable of accelerating drug discovery. To date, hypothesis h1. seems to be verified, (Nishikawa et al., 2014) while in the case of h3. we can partially rely on artificial neural networks and optimization techniques (Kingma & Ba, 2015; Graves, 2014).
@@
In this work, we take a step toward the realization of the joint technology conceptualized in h2. and h3. by proposing SynaptoGen [^1] https://github.com/BoCtrl-C (Figure 1), a model that links, by means of differentiable functions, vector representations of gene expression profiles (i.e., the pattern of genes actively being transcribed into RNA in a cell) and genetic rules (i.e., interaction probabilities of protein pairs involved in synaptogenesis) to the average number of synaptic connections between pairs of neurons as well as their synaptic conductance values. We substantiate our work through theoretical development, which hinges on novel propositions and related mathematical proofs. SynaptoGen is compatible with backpropagation and can be inserted in learning frameworks where optimization is performed through gradient descent, enabling management of network sizes and task complexities beyond the capabilities of other optimization techniques. Finally, SynaptoGen is designed with flexibility in mind, allowing practitioners to choose which biological quantities to optimize (e.g., genetic rules, expression profiles, or both).
@@
Figure 1: Multi-scale overview of the SynaptoGen model. A. In neural networks, synaptogenesis–the formation of synapses–can be approximated as the outcome of interactions (e.g., molecular binding) between proteins translated from gene pairs. SynaptoGen models this process as the realization $\tilde{B}_{uv}$ of a random variable defined as the sum of multiple binomial random variables (e.g., $\mathcal{B}^{AB}+\mathcal{B}^{BC}$), one for each potentially interacting gene pair. In the notation used, the $uv$ subscript indicates the neuronal pair taken into consideration. B. The use of binomial random variables stems from the idea of modeling synapse formation for each gene pair, as a process akin to flipping $n_{ij}$ biased coins, each representing a $\text{Bernoulli}(p_{ij})$ random variable. C. When working with matrix representations of the genetic factors in the panel, it is mathematically provable (see Section 2) that a series of matrix multiplications and point-wise operations yields a matrix where the entry in the $u$-th row and $v$-th column contains the expected number of synapses formed between neurons $u$ and $v$, multiplied by their average synaptic conductance. D. This resulting matrix, $W$ ($\bar{W}$ in Section 2), can be interpreted as a weight matrix and integrated into architectures such as multilayer perceptrons (MLPs). Here, $\bm{x}$ denotes an MLP layer’s input, while $\bm{a}$ the resulting activations.
@@
Related Work. Interest in leveraging organoid production for computational purposes has surged since 2022, following the introduction of DishBrain (Kagan et al., 2022) by Cortical Labs. DishBrain combines in-vitro neural networks, cultured from human or rodent sources, with a simulated environment–specifically the game ”Pong”–via a high-density multielectrode array. The system employs the free energy principle, which suggests that neural networks adapt by minimizing the unpredictability of their sensory inputs through belief updating and environmental interactions. Although the neuron cultures in DishBrain exhibited statistically significant improvements in gameplay performance (e.g., average rally length, % of aces and % of long-rallies) compared to the control conditions defined by the authors (i.e., random agents), it is difficult to conclude that the neuronal networks fully mastered the task.
@@
In contrast to DishBrain, our research aims to “train” networks of neurons by influencing synaptogenesis through genetic manipulations at the individual neuron level, a largely unexplored task in the literature, with only a limited number of closely related studies available. The basis for our work is outlined in (Barabási & Czégel, 2021; Barabási & Barabási, 2020), which introduces methods for constructing networks based on genetic encodings inspired by the wiring rules of the brain. These methods were further elaborated in the Connectome Model (CM) (Kovács et al., 2020), where the authors decomposed the adjacency matrix of a connectome into the product of three matrices representing specific genetic quantities. Another development was presented in (Barabási et al., 2023), where the CM’s matrix entries were treated as learnable parameters, resulting in the weight matrix of a Multilayer Perceptron (MLP) within the context of training neural networks. While this methodology has proven effective in producing parameter-efficient neural networks, it maintains a notable distance from the biological intricacies of real neuronal networks. A distinct generalization of the CM has also been proposed for the computational inference of synaptic polarities (Harris et al., 2022), a quantity not considered in the original CM.
@@
Similarly, our work draws inspiration from the CM but is geared towards a more bio-plausible computational modeling of synaptogenesis, with the novel elements extensively discussed in Section 2.
@@ 2 Methods
In the CM’s first interpretation, each row of $X$ referred to a specific neuron while the $i$-th entry of the row described the binary expression (1–“expressed”–or 0–“not expressed”) of gene $i$, one of the genes involved in synapse formation. Matrix $O$, instead, represented interaction compatibility, in the synapse formation process, for proteins translated from all gene pairs. Hence, $X\in\{0,1\}^{N\times G}$ and $O\in\{0,1\}^{G\times G}$ were defined as binary matrices while the entries of $B$, of shape $N\times N$, belonged to $\mathbb{Z^{+}}$; with $N$ and $G$ denoting the number of neurons and genes, respectively. When $G\ll N$ however, a very common scenario in nature (Koulakov et al., 2022), not all possible connectomes can be decomposed through equation 1. For this reason, the authors of the CM went on to relaxing the genetic rules matrix to $O\in[0,1]^{G\times G}$, interpreting its entries as probabilities, which generates the following approximation:
@@
$O$ $=\arg\min_{O^{\prime}}||B-XO^{\prime}X^{T}||^{2}$ (3)
where |||||$\bullet$||| | ∙ | | is intended as the Frobenius norm.
In this paper, we formulate a more general alternative to this framework by building a model that takes into account synaptic multiplicity and conductance. Starting from equation 1, we design two novel interpretations tightly linked to the formalism with which the quantities of interest (i.e., the number of synapses between neurons and their conductances) have been represented in our model. Our theoretical framework is as follows.
@@
In different terms, if the hypotheses of Proposition 1 are verified, gene expression in a pair of genes tells us how many attempts we can make to place a synapse between a pre- and a post-synaptic neuron; the genetic rule, instead, describes the probability of success, conditioned on the interaction between the proteins translated from the corresponding genes, for each attempt. It is worth noting that equation 6 represents the average number of links between two specific nodes of the connectome. Keeping in mind that genetic rules are shared across neurons, the equation can easily be generalized to the whole connectome:
@@
Let $\mathcal{T}$ be a multinomial random variable representing the process of randomly picking, from $u$, a synaptic vesicle filled with a specific neurotransmitter. And let $\mathcal{R}$ be a multinomial random variable representing the process of randomly selecting a specific receptor from the membrane of $v$. We define vectors $\bm{q}\in[0,1]^{L}$ and $\bm{r}\in[0,1]^{M}$ as the probability distributions associated with $\mathcal{T}$ and $\mathcal{R}$; where $L$ denotes the total number of neurotransmitters while $M$ the number of receptors. We also define $A\in\{-1,0,1\}^{L\times M}$ as the polarity matrix (for further details, refer to Appendix A) and $K\in{\mathbb{R}^{+}}^{L\times M}$ as the conductance matrix. In detail, entry $A_{ij}$ defines the polarity of synapses derived from the interaction of the $i$-th neurotransmitter with the $j$-th receptor ($A_{ij}=0$ if the considered neurotransmitter and receptor are not compatible) while $K_{ij}$ stores its linked conductance. We finally set $\mathcal{G}=f(\mathcal{T},\mathcal{R})$, with $f(i,j)=A_{ij}K_{ij}$. In other words, $\mathcal{G}$ represents the “signed” conductance of a synapse randomly selected from the ones that connect neurons $u$ and $v$.
@@
By expanding the matrix multiplications in equation 9, we have:
@@
As for equation 7, also equation 9 can be generalized by stacking the neurotransmitter distributions of all pre-synaptic neurons in $Q=[\dots,\bm{q},\dots]^{T}$ and the receptor distributions of post-synaptic neurons in $R=[\dots,\bm{r},\dots]^{T}$:
@@
As a next step, equation 7 and equation 10 can be inserted into the core equation of our model, which follows:
@@
Summarizing, through equation 11 we are able to express the average equivalent conductance between all pairs of neurons as a differentiable function of their gene expression profiles and distributions of synthesized neurotransmitters and receptors, which, in turn, depend on gene expression. Furthermore, in our formalism, synaptogenesis can be simulated by sampling from the random variables we have defined. For instance, the simplest approximation of synaptogenesis can be obtained as follows (further details in Appendices B, C):
@@ 3 Results
Our approach begins with a simplified setup, where a neuronal population comprises spatially-separated layers. We assumed that neurons in one layer could only attempt to form synapses with neurons in adjacent layers. This restriction allowed us to focus on multipartite network topologies and implement SynaptoGen as a customized MLP, where its weight matrices are decomposed according to equation 11, with genetic rules shared across layers.
@@
Initially, we separately trained SynaptoGen–16/32/64 genes, 3 neurotransmitters, neuronal population sizes of ${\sim}128$ (see Appendix D for the training details) on the four selected environments using the DQN algorithm (Mnih et al., 2015). The training provided the learned genetic rules, gene expression profiles, neurotransmitter and receptor distributions, and synaptic conductances that enabled the agents to perform the tasks. Notably, SynaptoGen is built on the average equivalent conductances introduced in equation 10, and can be interpreted as an average agent that reflects the effects of the underlying genetically-derived quantities.
Successively, for each environment, we sampled 100 neural networks from the trained SynaptoGen models, simulating the development, based on the computed synaptogenesis rules, of multiple populations of neurons into neuronal networks. We then measured their performance and compared the obtained metrics with those from two carefully designed baselines. The first baseline (Appendix E) adopts the optimization technique employed in (Stöckl et al., 2022), which has been applied to optimize related models, the probabilistic skeletons, for similar tasks, i.e., the optimization of connection probabilities between neuronal types. This technique, known as SNES (Schaul, 2012), is an evolutionary algorithm specifically designed for medium-to-large problem dimensions. Given the structure of SynaptoGen, which is explicitly designed for gradient-based optimization but remains general and flexible enough to accommodate other optimizers, SNES could serve as both a benchmark we aim to surpass and a potential alternative worth exploring. The second baseline (Appendix F) leveraged the SynaptoGen’s biology, using the optimized genetic rules $O$ and conductance matrix $K$, while initializing gene expression profiles in a bio-plausible manner following the procedure identified by Kerstjens et al. (Kerstjens et al., 2022; 2024). As before, 100 neuronal networks were sampled from each baseline.
@@
The results obtained from experiments with agents characterized by a genetic profile of 16 genes are shown in Figure 2. Table 1 provides summary metrics of the outcomes of each experiment. The first metric is the average reward calculated across groups of 100 sampled agents. The second metric is also the average reward, but calculated for the top-10 best-performing agents in each group. This metric aims to assess the models’ effectiveness in scenarios where the entity responsible for synthesizing biological agents prioritizes achieving a small subset of highly performing agents, even at the cost of the overall process yield–the number of agents capable of solving their task divided by the total number of agents. Finally, the third metric describes the percentage of simulated agents capable of solving the tasks defined by the four selected environments (see Appendix G for further discussion). This metric represents the global yield of each methodology under consideration.
In terms of average reward SynaptoGen (trained via gradient descent) emerged as the best-performing model, outperforming agents optimized with SNES in three out of four environments. It also consistently surpassed the baseline characterized by biologically plausible initialization of gene expression. The SNES agents also performed well, exceeding the bio-plausible baseline–the one characterized by the bio-plausible expression initialization–in three out of four environments. When examining the average rewards of the top-10 agents, the superiority of SynaptoGen and its derived agents becomes even more evident. They achieved the highest performance across all environments studied, sharing the leading position only in Cart Pole. By contrast, the SNES top-10 agents achieved average rewards indicative of task resolution in only 50% of the cases. For the bio-plausible baseline, nearly all top-10 agents failed to achieve sufficient performance. Finally, regarding the percentage of agents capable of solving the tasks their genetic profiles were optimized/initialized for, SynaptoGen was the only approach to achieve non-zero percentages in all four environments. Moreover, it achieved the highest percentage in three out of four cases. SNES agents also performed reasonably well, reaching a 100% resolution rate in half the environments, though failing entirely in the other half. For the bio-plausible baseline, only 2 out of 400 sampled agents successfully solved their respective tasks. Notably, in 10 out of 12 comparisons, the differences between reward distributions were statistically significant based on the Mann-Whitney test with Bonferroni correction (for the multiple environment-specific comparisons performed). For brevity and clarity, results for agents derived from genetic profiles with 32 or 64 genes are presented in Appendix H. These results exhibit minimal differences from those reported here and lead to equivalent conclusions.
@@
To provide additional evidence of SynaptoGen’s effectiveness, we designed an additional set of simulations aimed at validating the framework in a context closer to the biological reality of neuronal networks. In these experiments, we used genetic rules derived from experimental data collected on the nerve ring of the nematode C. elegans. This region was selected due to the availability of both transcriptomic data (Taylor et al., 2021) and connectomic data–including membrane contacts and synapses (Brittin et al., 2021; Cook et al., 2019; Witvliet et al., 2020)–collected from it. The genetic rules were constructed based on a restricted set of 141 cell adhesion molecules (CAMs) with a documented role in synapse formation (Colón-Ramos et al., 2007; Kim & Emmons, 2017), following a procedure adapted from the Network Differential Gene Expression Analysis (nDGE) framework proposed in (Taylor et al., 2021). The procedure is described in Appendix I and summarized in Figure 3. Briefly, we first identified pairs of genes responsible for CAM production that were co-expressed in nerve ring neurons with synaptic connections but not co-expressed in cells sufficiently close to form synapses yet lacking connections. We then constrained the (learnable) genetic linked to these pairs to assume high probabilities ($>0.5$), while assigning low probabilities ($<0.5$) to all other rules.

### v3

@@ Abstract
There is growing consensus among neuroscientists that neural circuits critical for survival are the result of genomic decompression processes. We introduce SynaptoGen, a novel computational framework–member of the Connectome Models family–bringing synthetic biological intelligence closer, facilitating neural biological agent development through precise genetic control of synaptogenesis. SynaptoGen is the first model of its kind offering mechanistic explanation of synaptic multiplicity based on genetic expression and protein interaction probabilities. The framework connects genetic factors through a differentiable function, working as a neural network where synaptic weights equal average numbers of synapses between neurons, multiplied by conductance, derived from genetic profiles. Differentiability enables gradient-based optimization, allowing generation of genetic expression patterns producing pre-wired biological agents for specific tasks. Validation in simulated synaptogenesis scenarios shows agents successfully solving four reinforcement learning benchmarks, consistently surpassing control baselines. Despite gaps in biological realism requiring mitigation, this framework has potential to accelerate synthetic biological intelligence research.
@@ 1 Introduction
Let us consider a scenario in which, during brain development, hypothesis 1 - h1. it becomes possible to manipulate, prior to the formation of synapses, the gene expression profiles of individual neurons. Furthermore, let us assume that h2. we possess the capability to influence these expression profiles in a manner that directs synaptogenesis towards a specific neuronal network topology. In this context, h3. we might be able to obtain the optimal computational graph, expressed as a composition of functions that represent the behaviour of neurons, required to solve a task of interest. Small living organisms, or organoids (Bhaduri et al., 2020), could be, in principle, genetically programmed to develop into neuronal networks capable of solving pre-specified tasks. Such technology would lead to disruptive applications–e.g., extreme low-power computing, micro-devices for the control of biological systems, or novel biological testing platforms capable of accelerating drug discovery. To date, hypothesis h1. seems to be verified, (Nishikawa et al., 2014) while in the case of h3. we can partially rely on artificial neural networks and optimization techniques (Kingma and Ba, 2015; Graves, 2014).
@@
In this work, we take a step toward the realization of the joint technology conceptualized in h2. and h3. by proposing SynaptoGen [^1] https://github.com/BoCtrl-C/synaptogen (Figure 1), a model that links, by means of differentiable functions, vector representations of gene expression profiles (i.e., the pattern of genes actively being transcribed into RNA in a cell) and genetic rules (i.e., interaction probabilities of protein pairs involved in synaptogenesis) to the average number of synaptic connections between pairs of neurons as well as their synaptic conductance values. We substantiate our work through theoretical development, which hinges on novel propositions and related mathematical proofs. SynaptoGen is compatible with backpropagation and can be inserted in learning frameworks where optimization is performed through gradient descent, enabling management of network sizes and task complexities beyond the capabilities of other optimization techniques. Finally, SynaptoGen is designed with flexibility in mind, allowing practitioners to choose which biological quantities to optimize (e.g., genetic rules, expression profiles, or both).
@@
Figure 1: Multi-scale overview of the SynaptoGen model. A. In neural networks, synaptogenesis–the formation of synapses–can be approximated as the outcome of interactions (e.g., molecular binding) between proteins translated from gene pairs. SynaptoGen models this process as the realization $\tilde{B}_{uv}$ of a random variable defined as the sum of multiple binomial random variables (e.g., $\mathcal{B}^{AB}+\mathcal{B}^{BC}$), one for each potentially interacting gene pair. In the notation used, the uvuv subscript indicates the neuronal pair taken into consideration. B. The use of binomial random variables stems from the idea of modeling synapse formation for each gene pair, as a process akin to flipping $n_{ij}$ biased coins, each representing a $\text{Bernoulli}(p_{ij})$ random variable. C. When working with matrix representations of the genetic factors in the panel, it is mathematically provable (see Section Methods) that a series of matrix multiplications and point-wise operations yields a matrix where the entry in the $u$-th row and $v$-th column contains the expected number of synapses formed between neurons $u$ and $v$, multiplied by their average synaptic conductance. D. This resulting matrix, $W$ ($\bar{W}$ in Section Methods), can be interpreted as a weight matrix and integrated into architectures such as multilayer perceptrons (MLPs). Here, $\bm{x}$ denotes an MLP layer’s input, while $\bm{a}$ the resulting activations.
1.1 Related Work
@@
Interest in leveraging organoid production for computational purposes has surged since 2022, following the introduction of DishBrain (Kagan et al., 2022) by Cortical Labs. DishBrain combines in-vitro neural networks, cultured from human or rodent sources, with a simulated environment–specifically the game "Pong"–via a high-density multielectrode array. The system employs the free energy principle, which suggests that neural networks adapt by minimizing the unpredictability of their sensory inputs through belief updating and environmental interactions. Although the neuron cultures in DishBrain exhibited statistically significant improvements in gameplay performance (e.g., average rally length, % of aces and % of long-rallies) compared to the control conditions defined by the authors (i.e., random agents), it is difficult to conclude that the neuronal networks fully mastered the task.
@@
In contrast to DishBrain, our research aims to “train” networks of neurons by influencing synaptogenesis through genetic manipulations at the individual neuron level, a largely unexplored task in the literature, with only a limited number of closely related studies available. The basis for our work is outlined in (Barabási and Czégel, 2021; Barabási and Barabási, 2020), which introduces methods for constructing networks based on genetic encodings inspired by the wiring rules of the brain. These methods were further elaborated in the Connectome Model (CM) (Kovács et al., 2020), where the authors decomposed the adjacency matrix of a connectome into the product of three matrices representing specific genetic quantities. Another development was presented in (Barabási et al., 2023), where the CM’s matrix entries were treated as learnable parameters, resulting in the weight matrix of a Multilayer Perceptron (MLP) within the context of training neural networks. While this methodology has proven effective in producing parameter-efficient neural networks, it maintains a notable distance from the biological intricacies of real neuronal networks. A distinct generalization of the CM has also been proposed for the computational inference of synaptic polarities (Harris et al., 2022), a quantity not considered in the original CM.
@@
Similarly, our work draws inspiration from the CM but is geared towards a more bio-plausible computational modeling of synaptogenesis, with the novel elements extensively discussed in Section Methods.
@@ 2 Methods
In the CM’s first interpretation, each row of $X$ referred to a specific neuron while the $i$-th entry of the row described the binary expression (1–“expressed”–or 0–“not expressed”) of gene $i$, one of the genes involved in synapse formation. Matrix $O$, instead, represented interaction compatibility, in the synapse formation process, for proteins translated from all gene pairs. Hence, $X\in\{0,1\}^{N\times G}$ and $O\in\{0,1\}^{G\times G}$ were defined as binary matrices while the entries of $B$, of shape $N\times N$, belonged to $\mathbb{Z^{+}}$; with $N$ and $G$ denoting the number of neurons and genes, respectively. When $G\ll N$ however, a very common scenario in nature (Koulakov et al., 2022), not all possible connectomes can be decomposed through (1). For this reason, the authors of the CM went on to relaxing the genetic rules matrix to $O\in[0,1]^{G\times G}$, interpreting its entries as probabilities, which generates the following approximation:
@@
[…]
where $||\bullet||$ is intended as the Frobenius norm.
In this paper, we formulate a more general alternative to this framework by building a model that takes into account synaptic multiplicity and conductance. Starting from (1), we design two novel interpretations tightly linked to the formalism with which the quantities of interest (i.e., the number of synapses between neurons and their conductances) have been represented in our model. Our theoretical framework is as follows.
@@
In different terms, if the hypotheses of Proposition 1 are verified, gene expression in a pair of genes tells us how many attempts we can make to place a synapse between a pre- and a post-synaptic neuron; the genetic rule, instead, describes the probability of success, conditioned on the interaction between the proteins translated from the corresponding genes, for each attempt. It is worth noting that (6) represents the average number of links between two specific nodes of the connectome. Keeping in mind that genetic rules are shared across neurons, the equation can easily be generalized to the whole connectome:
@@
Let $\mathcal{T}$ be a multinomial random variable representing the process of randomly picking, from $u$, a synaptic vesicle filled with a specific neurotransmitter. And let $\mathcal{R}$ be a multinomial random variable representing the process of randomly selecting a specific receptor from the membrane of $v$. We define vectors $\bm{q}\in[0,1]^{L}$ and $\bm{r}\in[0,1]^{M}$ as the probability distributions associated with $\mathcal{T}$ and $\mathcal{R}$; where $L$ denotes the total number of neurotransmitters while $M$ the number of receptors. We also define $A\in\{-1,0,1\}^{L\times M}$ as the polarity matrix (for further details, refer to Appendix The Polarity Matrix) and $K\in{\mathbb{R}^{+}}^{L\times M}$ as the conductance matrix. In detail, entry $A_{ij}$ defines the polarity of synapses derived from the interaction of the $i$-th neurotransmitter with the $j$-th receptor ($A_{ij}=0$ if the considered neurotransmitter and receptor are not compatible) while $K_{ij}$ stores its linked conductance. We finally set $\mathcal{G}=f(\mathcal{T},\mathcal{R})$, with $f(i,j)=A_{ij}K_{ij}$. In other words, $\mathcal{G}$ represents the “signed” conductance of a synapse randomly selected from the ones that connect neurons $u$ and $v$.
@@
By expanding the matrix multiplications in (9), we have:
@@
As for (7), also (9) can be generalized by stacking the neurotransmitter distributions of all pre-synaptic neurons in $Q=[\dots,\bm{q},\dots]^{T}$ and the receptor distributions of post-synaptic neurons in $R=[\dots,\bm{r},\dots]^{T}$:
@@
As a next step, (7) and (10) can be inserted into the core equation of our model, which follows:
@@
Summarizing, through (11) we are able to express the average equivalent conductance between all pairs of neurons as a differentiable function of their gene expression profiles and distributions of synthesized neurotransmitters and receptors, which, in turn, depend on gene expression. Furthermore, in our formalism, synaptogenesis can be simulated by sampling from the random variables we have defined. For instance, the simplest approximation of synaptogenesis can be obtained as follows (further details in Appendices Simulating Synaptogenesis, A Problem of Quantization):
@@ 3 Results
Our approach begins with a simplified setup, where a neuronal population comprises spatially-separated layers. We assumed that neurons in one layer could only attempt to form synapses with neurons in adjacent layers. This restriction allowed us to focus on multipartite network topologies and implement SynaptoGen as a customized MLP, where its weight matrices are decomposed according to (11), with genetic rules shared across layers.
@@
Initially, we separately trained SynaptoGen–16/32/64 genes, 3 neurotransmitters, neuronal population sizes of ${\sim}128$ (see Appendix Training Details for the training details)–on the four selected environments using the DQN algorithm (Mnih et al., 2015). The training provided the learned genetic rules, gene expression profiles, neurotransmitter and receptor distributions, and synaptic conductances that enabled the agents to perform the tasks. Notably, SynaptoGen is built on the average equivalent conductances introduced in (10), and can be interpreted as an average agent that reflects the effects of the underlying genetically-derived quantities.
Successively, for each environment, we sampled 100 neural networks from the trained SynaptoGen models, simulating the development, based on the computed synaptogenesis rules, of multiple populations of neurons into neuronal networks. We then measured their performance and compared the obtained metrics with those from two carefully designed baselines. The first baseline (Appendix The Separable Natural Evolution Strategy (SNES) Baseline) adopts the optimization technique employed in (Stöckl et al., 2022), which has been applied to optimize related models, the probabilistic skeletons, for similar tasks, i.e., the optimization of connection probabilities between neuronal types. This technique, known as SNES (Schaul, 2012), is an evolutionary algorithm specifically designed for medium-to-large problem dimensions. Given the structure of SynaptoGen, which is explicitly designed for gradient-based optimization but remains general and flexible enough to accommodate other optimizers, SNES could serve as both a benchmark we aim to surpass and a potential alternative worth exploring. The second baseline (Appendix The Bio-Plausible Baseline) leveraged the SynaptoGen’s biology, using the optimized genetic rules $O$ and conductance matrix $K$, while initializing gene expression profiles in a bio-plausible manner following the procedure identified by Kerstjens et al. (Kerstjens et al., 2022, 2024). As before, 100 neuronal networks were sampled from each baseline.
@@
The results obtained from experiments with agents characterized by a genetic profile of 16 genes are shown in Figure 2. Table 1 provides summary metrics of the outcomes of each experiment. The first metric is the average reward calculated across groups of 100 sampled agents. The second metric is also the average reward, but calculated for the top-10 best-performing agents in each group. This metric aims to assess the models’ effectiveness in scenarios where the entity responsible for synthesizing biological agents prioritizes achieving a small subset of highly performing agents, even at the cost of the overall process yield–the number of agents capable of solving their task divided by the total number of agents. Finally, the third metric describes the percentage of simulated agents capable of solving the tasks defined by the four selected environments (see Appendix Reward Thresholds for Considering a Task Solved for further discussion). This metric represents the global yield of each methodology under consideration.
In terms of average reward SynaptoGen (trained via gradient descent) emerged as the best-performing model, outperforming agents optimized with SNES in three out of four environments. It also consistently surpassed the baseline characterized by biologically plausible initialization of gene expression. The SNES agents also performed well, exceeding the bio-plausible baseline–the one characterized by the bio-plausible expression initialization–in three out of four environments. When examining the average rewards of the top-10 agents, the superiority of SynaptoGen and its derived agents becomes even more evident. They achieved the highest performance across all environments studied, sharing the leading position only in Cart Pole. By contrast, the SNES top-10 agents achieved average rewards indicative of task resolution in only 50% of the cases. For the bio-plausible baseline, nearly all top-10 agents failed to achieve sufficient performance. Finally, regarding the percentage of agents capable of solving the tasks their genetic profiles were optimized/initialized for, SynaptoGen was the only approach to achieve non-zero percentages in all four environments. Moreover, it achieved the highest percentage in three out of four cases. SNES agents also performed reasonably well, reaching a 100% resolution rate in half the environments, though failing entirely in the other half. For the bio-plausible baseline, only 2 out of 400 sampled agents successfully solved their respective tasks. Notably, in 10 out of 12 comparisons, the differences between reward distributions were statistically significant based on the Mann-Whitney test with Bonferroni correction (for the multiple environment-specific comparisons performed). For brevity and clarity, results for agents derived from genetic profiles with 32 or 64 genes are presented in Appendix Additional Results. These results exhibit minimal differences from those reported here and lead to equivalent conclusions.
@@
To provide additional evidence of SynaptoGen’s effectiveness, we designed an additional set of simulations aimed at validating the framework in a context closer to the biological reality of neuronal networks. In these experiments, we used genetic rules derived from experimental data collected on the nerve ring of the nematode C. elegans. This region was selected due to the availability of both transcriptomic data (Taylor et al., 2021) and connectomic data–including membrane contacts and synapses (Brittin et al., 2021; Cook et al., 2019; Witvliet et al., 2020)–collected from it. The genetic rules were constructed based on a restricted set of 141 cell adhesion molecules (CAMs) with a documented role in synapse formation (Colón-Ramos et al., 2007; Kim and Emmons, 2017), following a procedure adapted from the Network Differential Gene Expression Analysis (nDGE) framework proposed in (Taylor et al., 2021). The procedure is described in Appendix nDGE-based Bio-Plausible Rules and summarized in Figure 3. Briefly, we first identified pairs of genes responsible for CAM production that were co-expressed in nerve ring neurons with synaptic connections but not co-expressed in cells sufficiently close to form synapses yet lacking connections. We then constrained the (learnable) genetic linked to these pairs to assume high probabilities (>0.5>0.5), while assigning low probabilities (<0.5<0.5) to all other rules.